OpenStudy (mikeymaximum):

What is 5/6 divided by 1/3 A.5/18 B.8/14 C.7 1/2 D.2 1/2

OpenStudy (alexandervonhumboldt2):

5/6*3=15/6=2 1/2

OpenStudy (mikeymaximum):

Thx